### v2

# Right Splitting, Galois Correspondence, Galois Representations and Inverse Galois Problem

Abstract.
In this article, we realize some groups as Galois groups over rational numbers and finite extension of rational numbers by studying right splitting of some exact sequences, Galois correspondence and algebraic operations on Galois representations.
2020 Mathematics Subject Classification:
11R32, 12F12, 11F80
1. Introduction
The famous ‘Inverse Galois problem’ asks whether every finite group appears as the Galois group of some finite Galois extension of $\mathbb{Q}$}. Many families of simple groups are known instances of this problem but the general question is still open.
Using Galois representations attached to elliptic curves, Arias-de-Reyna and König have proved in [1] that there are infinitely many number fields with Galois group ${\rm GL}_{2}({\mathbb{F}}_{p})$ over $\mathbb{Q}$} that are pairwise linearly disjoint over ${\mathbb{Q}}(\sqrt{p^{\ast}})$ (see Remark 2.6.2).
In [8], the case of occurrence of ${\rm PSL}_{2}({\mathbb{F}}_{p})$ as a Galois group over $\mathbb{Q}$} for all primes $p\geq 5$ was established by Zywina using the results of Ribet [7] about the Deligne’s Galois representations associated to certain newforms.
In this article, we establish some instances of Inverse Galois problem. In Sec. 2, through Galois correspondence and right splitting of some exact sequences of groups, we obtain the main results Theorem 2.4, Theorem 2.6. We then apply these general results to the case in [1] and obtain interesting consequences as corollaries which are as follows.
Theorem 1.1.
Let $p$ be a prime $\geq 5$.
  1. (1)
    The group $({\rm SL}_{2}(\mathbb{F}_{p})\times{\rm SL}_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/(p-1)\mathbb{Z}$ (with semidirect product group law as in Thm 2.6) is realizable as Galois group over $\mathbb{Q}$.
  2. (2)
    The group $({\rm PSL}_{2}(\mathbb{F}_{p})\times{\rm PSL}_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/2\mathbb{Z}$ (with semidirect product group law as in Thm 2.4) is realizable as Galois group over $\mathbb{Q}$.
  3. (3)
    For the case when $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ (hence normal) subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{p})/H$ (with semidirect product group law as in Thm 2.6) is realizable as Galois group over $\mathbb{Q}$.
In Sec. 3, using the algebraic operation induction on Galois representations and right splitting of some exact sequences of groups, we obtain Thm. 3.7, Thm. 3.8 as the main results. We then apply these general results to the case in [8] and obtain following interesting consequences in Thm. 3.12 and Thm. 3.13.
Theorem 1.2.
  1. (1)
    For a prime $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realizable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Thm 3.7). The group obtained here is not isomorphic to ${\rm PGL}_{2}(\mathbb{F}_{p})$.
  2. (2)
    The group ${\rm PSL}_{2}(\mathbb{F}_{q})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7) for (1) $q=p$ for $p\geq 5$, (2) $q=p^{2}$ for $p\equiv\pm 2\ mod\ 5,\ p\geq 7$, (3) $q=p^{2}$ for $p\equiv\pm 3\ mod\ 8,\ p\geq 5$, (4) $q=p^{3}$ for odd prime $p\equiv\pm 2,\pm 3,\pm 4,\pm 6\ mod\ 13$ and (5) $q=5^{3},3^{5},3^{4}$.
Then for the algebraic operations direct sum and tensor products on Galois representations we obtain an unanticipated and interesting result Prop 3.15.
Then by using the algebraic operations direct sum and tensor products on Galois representations and right splitting of some exact sequences of groups, we obtain Thm. 3.19 as the main result. We then apply this general result to the case in [8] and obtain that ${\rm PSL}_{2}(\mathbb{F}_{p})\times\mathbb{Z}/2\mathbb{Z}$ is realizable as Galois Group over $\mathbb{Q}$ for $p\geq 5$.
Right Splitting of exact sequences of groups is the common thread that runs through the whole article.
2. Groups, Galois Correspondence, Right Splitting and Inverse Galois Problem
In this section, we describe some group theoretic results and then through these results as well as Galois correspondence and right splitting of some exact sequences, we obtain some general Galois theoretic results. Then we apply these general results to the cases described in [1] and obtain interesting consequences. Although some of the results related to groups and Galois correspondence may be known to experts, but we include them in our discussion for the sake of completeness.
2.1. Right splitting and group theoretic results
Recall that if $H$ is a normal subgroup of $G$ such that $G/H$ is abelian then $G,G]\subset H$. Following is a simple lemma where the converse follows from structure theorem for abelian groups.
Lemma 2.1.
Let $G$ be a finite group. Then the condition $G/[G,G]$ is cyclic of order $m$ is equivalent to the condition that for any $n|m$, $G$ has a unique index-$n$ normal subgroup such that the quotient is abelian (the quotient with that subgroup is in fact cyclic). This unique subgroup is given by
$H=\{x\in G\ |\ \pi(x)\text{ is an}\ n\text{-th power in}\ G/[G,G]\}$
where $\pi$ is the quotient homomorphism $G\rightarrow G/[G,G]$.
Let $\mathbb{F}_{q}$ be finite field of $q$ elements where $q=p^{r}$ for an odd prime $p$ and $r\in\mathbb{N}$. We know that ${\rm GL}_{2}(\mathbb{F}_{q}),{\rm GL}_{2}(\mathbb{F}_{q})]=SL_{2}(\mathbb{F}_{q})$ and ${\rm GL}_{2}(\mathbb{F}_{q})/SL_{2}(\mathbb{F}_{q})\cong\mathbb{F}_{q}^{\times}$ through the surjective determinant map and $\mathbb{F}_{q}^{\times}$ is cyclic. By applying similar argument to detdet map as applied to $\pi$ map in above lemma, we get the following.
Corollary 2.1.1.
Let $n\in\mathbb{N}$. If $q\equiv 1$ modmod $n$, then ${\rm GL}_{2}(\mathbb{F}_{q})$ has a unique index-$n$ normal subgroup such that the quotient is abelian. (The quotient with that subgroup is in fact cyclic). This unique subgroup is given by
$H=\{x\in{\rm GL}_{2}(\mathbb{F}_{q})\ |\ det(x)\ \text{ is an}\ n\text{-th power in }\ \mathbb{F}_{q}^{\times}\}.$
Corollary 2.1.2.
  1. (1)
    Let $G$ be a group. If $[G,G]$ has index $m$ in $G$, then $[G,G]$ is the unique index-$m$ subgroup of $G$ such that the quotient is abelian.
  2. (2)
    $SL_{2}(\mathbb{F}_{q})$ is the unique index-(q1)(q-1) subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
For $a,b\in\mathbb{N}$, we denote $gcd(a,b)$ by $(a,b)$.
Proposition 2.2.
Let $n\in\mathbb{N}$ and let $G$ be a group such that $G/[G,G]$ is cyclic of order $m$. Let $n|m$ and let $H$ be the unique index-$n$ normal subgroup of $G$ such that the quotient is abelian. Then the following holds.
  1. (1)
    If there exists $x\in G$ such that $x^{n}=1$ and $\{1,x,x^{2},...,x^{n-1}\}$ is a set of representatives for $H$-cosets in $G$, then (n,m/n) =1.
  2. (2)
    If $(n,m/n)=1$ and the exact sequence
    (2.1) $1\rightarrow[G,G]\rightarrow G\rightarrow G/[G,G]\rightarrow 1$
    is right split, then there exists $x\in G$ such that $x^{n}=1$ and $\{1,x,x^{2},...,x^{n-1}\}$ is a set of representatives for $H$-cosets in $G$.
  3. (3)
    Let $n=m$ (hence $[G,G]$ is the unique index-$m$ subgroup of $G$ such that quotient is abelian). Then the exact sequence 2.1 is right split if and only if there exists $x\in G$ such that $x^{m}=1$ and $\{1,x,x^{2},...,x^{m-1}\}$ is a set of representatives for $[G,G]$-cosets in $G$.
Proof.
  1. (1)
    Suppose we have a set of representatives for $H$-cosets in $G$ of the form $\{1,x,x^{2},...,x^{n-1}\}$ with $x\in G$ such that $x^{n}=1$. Let $y$ be a generator of cyclic group $G/[G,G]$ and let $0\leq l\leq(m-1)$ be such that $\pi(x)=y^{l}$. As $x^{n}=1$, we have $1=\pi(x)^{n}=y^{ln}$ and hence $m\mid ln$. Thus we have j = ln/m ∈Z and 1≤j≤(n-1).
    Now assume on the contrary that $(n,m/n)\neq 1$. Then for $k=n/(n,m/n)\leq(n-1)$, we have $n\mid km/n$, hence $n\mid kl$, and therefore $\pi(x^{k})=y^{lk}$ is an $n$-th power in $G/[G,G]$, hence $x^{k}\in H$. Hence we get a contradiction. Similarly we can show that $(j,n)=1$.
  2. (2)
    If $(n,m/n)=1$, then we clearly have that $n\nmid km/n$ for all $1\leq k\leq(n-1)$. Consider the quotient map $\pi:G\rightarrow G/[G,G]$. Since given exact sequence 2.1 is right split, we have an injective homomorphism $\iota:G/[G,G]\rightarrow G$ such that $\pi\circ\iota=id_{G/[G,G]}$. Let $x=\iota(y^{jm/n})\in G$ where $1\leq j\leq(n-1)$ and $(j,n)=1$. Hence $\pi(x)=y^{jm/n}$. Since $\iota$ is injective, $x$ will be an element of order $n$. Now since $n\nmid km/n$ for all $1\leq k\leq(n-1)$ and since $(j,n)=1$, we have $n\nmid jkm/n$ for all $1\leq k\leq(n-1)$. Hence for all $1\leq k\leq(n-1)$, $\pi(x^{k})=y^{jkm/n}$ is not an $n$-th power in $G/[G,G]$. Hence $x^{k}\not\in H$. Therefore we get the required set of representatives for $H$-cosets in $G$.
  3. (3)
    If the exact sequence 2.1 is right split, then by (2), we are done. The converse is easy.
Remark 2.2.1.
  1. (1)
    If $G$ is abelian, then it satisfies the hypothesis of above proposition if and only if it is cyclic.
  2. (2)
    The condition that $G/[G,G]$ is cyclic is important in statement (3) above. If $G$ is a non-cyclic abelian group, then $[G,G]=1$ and so exact sequence $1\rightarrow 1\rightarrow G\rightarrow G\rightarrow 1$ is clearly right split, but we can never have required representatives for $[G,G]$-cosets because of non-cyclicity of $G$.
Corollary 2.2.1.
Let $n\in\mathbb{N}$. Consider $q$ such that $q\equiv 1$ mod $n$. Let $H$ be the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian. Then the following are equivalent.
  • (n,(q1)/n)=1(n,(q-1)/n)=1.
  • There exists a set of representatives for $H$-cosets in ${\rm GL}_{2}(\mathbb{F}_{q})$ of the form $\{1,x,x^{2},...,x^{n-1}\}$ with $x\in{\rm GL}_{2}(\mathbb{F}_{q})$ such that $x^{n}=I$.
Proof.
Let $G={\rm GL}_{2}(\mathbb{F}_{q})$. We know that below exact sequence given by the determinant map is right split.
(2.2) $1\rightarrow[G,G]\rightarrow G\xrightarrow{det}\mathbb{F}_{q}^{\times}\rightarrow 1.$
By applying similar argument to detdet map as applied to $\pi$ map in previous proposition, we are done.∎
The following follows from Dirichlet’s theorem on primes in arithmetic progression.
Proposition 2.3.
For given $n,r\in\mathbb{N}$ such that $(r,n)=1$, there are infinitely many primes satisfying the two conditions
$p^{r}\equiv 1\ mod\ n\ \text{and}\ (n,(p^{r}-1)/n)=1.$
2.2. Galois correspondence, right splitting and Galois theoretic results
We discuss a result here that establishes the occurrence of certain semidirect product of groups as a Galois group.
Theorem 2.4.
Let $l\geq 2$. Let $K$ be a finite extension of $\mathbb{Q}$ and let $E_{1},E_{2},...,E_{l}$ be finite Galois extensions of $\mathbb{Q}$ contained in $\bar{\mathbb{Q}}$. Let $G_{i}$ be the Galois group of $E_{i}$ over $\mathbb{Q}$ for all $i$. Suppose
  1. (1)
    $E_{1}\cdots E_{k}\cap E_{k+1}=K$ for all $1\leq k\leq l-1$,
  2. (2)
    for every $i$, there exists a set of representatives of $H_{i}$-cosets in respective $G_{i}$ that is closed under multiplication, where $H_{i}$ is Galois group of $E_{i}$ over $K$.
Then for each $1\leq i\leq l$, $H_{i}$ is a normal subgroup of respective $G_{i}$ and the group $(H_{1}\times H_{2}\times\dots\times H_{l})\rtimes G_{1}/H_{1}$ (for some semidirect product group law) is realizable as the Galois Group of $E_{1}\cdots E_{l}$ over $\mathbb{Q}$.
Proof.
We will prove by induction that ${\rm Gal}(E_{1}E_{2}\cdots E_{k}/K)\cong(H_{1}\times H_{2}\times\dots\times H_{k})$ and ${\rm Gal}(E_{1}E_{2}\cdots E_{k}/\mathbb{Q})\cong(H_{1}\times H_{2}\times\dots\times H_{k})\rtimes G_{1}/H_{1}$ (for some semidirect product group law) for all $2\leq k\leq l$.
For base case $k=2$, see the following diagram.
Since, $E_{1},E_{2}$ are Galois over $\mathbb{Q}$ and are contained in $\bar{\mathbb{Q}}$, $E_{1}E_{2}$ and $E_{1}\cap E_{2}$ are defined and are Galois over $Q$ and $K=E_{1}\cap E_{2}$. We have ${\rm Gal}(E_{i}/\mathbb{Q})=G_{i}$ and ${\rm Gal}(E_{i}/K)=H_{i}$ for $i=1,2$. Since $K$ is Galois over $\mathbb{Q}$, $H_{i}$s are normal in respective $G_{i}$s. ${\rm Gal}(K/\mathbb{Q})\cong G_{1}/H_{1}\cong G_{2}/H_{2}$. Fix an isomorphism $\theta$ from $G_{2}/H_{2}$ to $G_{1}/H_{1}$.
$\mathbb{Q}$$E_{1}\cap E_{2}$$E_{2}$$E_{1}E_{2}$$E_{1}$$H_{2}$$H_{1}$$H_{2}$$H_{1}$$H_{1}\times H_{2}$$G_{1}/H_{1}$$G_{1}$$G_{2}$$M$
Let ${\rm Gal}(K/\mathbb{Q})=\{x_{i}H_{1}|1\leq i\leq n\}$ where $x_{i}H_{1}=\theta(y_{i}H_{2})$ for all $i$. Here $\{x_{i}\in G_{1}|1\leq i\leq n\}$ and $\{y_{i}\in G_{2}|1\leq i\leq n\}$ are sets of representatives of $H_{1}$ and $H_{2}$ cosets in $G_{1}$ and $G_{2}$ respectively, which are closed under multiplication.
By Galois Correspondence [2, Theorem 2.1] and [2, Theorem 2.6], we have ${\rm Gal}(E_{1}E_{2}/E_{1})\cong H_{2}$, ${\rm Gal}(E_{1}E_{2}/E_{2})\cong H_{1}$ and ${\rm Gal}(E_{1}E_{2}/K)\cong H_{1}\times H_{2}$.
Let ${\rm Gal}(E_{1}E_{2}/\mathbb{Q})=M$. We have
$\begin{split}|M|&=[E_{1}E_{2}:K]\ [K:\mathbb{Q}]=|H_{1}\times H_{2}|\ |G_{1}/H_{1}|\\ &=|H_{1}\times H_{2}|\ |G_{2}/H_{2}|=|G_{1}|\ |H_{2}|=|G_{2}|\ |H_{1}|.\end{split}$
Consider map $\sigma:M\rightarrow G_{1}\times G_{2}$ given by $\sigma(g)=(g_{1},g_{2})$, where $g_{i}=g|_{E_{i}}$ for $i=1,2$. The map $\sigma$ is clearly a well defined injective group homomorphism (see [2, Theorem 1.1]).
If $g\in M$ and $g|_{K}=x_{i}H_{1}=g_{1}|_{K}$, then $(x_{i}^{-1}g_{1})|_{K}=id_{K}$. Hence $x_{i}^{-1}g_{1}\in H_{1}$, thus $g_{1}\in x_{i}H_{1}$. Similarly, by $g|_{K}=x_{i}H_{1}=\theta(g_{2}|_{K})=\theta(y_{i}H_{2})$, $g_{2}\in y_{i}H_{2}$.
Hence ${\rm Image}(\sigma)\subset\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})$. Since $|\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})|=|H_{1}\times H_{2}||G_{1}/H_{1}|=|M|$ and $\sigma$ is injective, we get ${\rm Image}(\sigma)=\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})$. Hence, $M\cong\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})$.
Consider a map $\psi:\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})\rightarrow(H_{1}\times H_{2})\rtimes G_{1}/H_{1}$ given by $\psi((x_{i}h_{1},y_{i}h_{2}))=((h_{1},h_{2}),x_{i}H_{1})$ where we define a semidirect product group law for $(H_{1}\times H_{2})\rtimes G_{1}/H_{1}$ by
$((h_{1},h_{2}),x_{i}H_{1}).((k_{1},k_{2}),x_{j}H_{1})=((x_{j}^{-1}h_{1}x_{j}k_{1},y_{j}^{-1}h_{2}y_{j}k_{2}),x_{i}x_{j}H_{1}).$
This group law is well defined and associative since $H_{i}$ are normal in $G_{i}$ for $i=1,2$ and sets of their respective coset representatives are closed under multiplication. We observe that
$(x_{i}h_{1},y_{i}h_{2}).(x_{j}k_{1},y_{j}k_{2})=(x_{i}h_{1}x_{j}k_{1},y_{i}h_{2}y_{j}k_{2})=((x_{i}x_{j})(x_{j}^{-1}h_{1}x_{j}k_{1}),(y_{i}y_{j})(y_{j}^{-1}h_{2}y_{j}k_{2})).$
Hence we conclude that $\psi$ is a group isomorphism and hence $M\cong(H_{1}\times H_{2})\rtimes G_{1}/H_{1}$.
Alternatively, consider the sequence
(2.3) $1\rightarrow(H_{1}\times H_{2})\xrightarrow{i}\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})\xrightarrow{\pi}G_{1}/H_{1}\rightarrow 1$
where $i(h_{1},h_{2})=(x_{1}h_{1},y_{1}h_{2})$ ($x_{1}\in H_{1},y_{1}\in H_{2}$) and $\pi(x_{i}h_{1},y_{i}h_{2})=x_{i}H_{1}$. Because of our multiplicatively closed assumption, $x_{1}$ and $y_{1}$ are identities of $G_{1}$ and $G_{2}$ respectively. Hence $i$ is injective. Also $\pi$ is surjective and $ker(\pi)=Image(i)$. Thus the sequence is exact.
Now consider $\iota:G_{1}/H_{1}\rightarrow\bigsqcup\limits_{i=1}^{n}(x_{i}H_{1}\times y_{i}H_{2})$ with $\iota(x_{i}H_{1})=(x_{i},y_{i})$. The map $\iota$ is clearly a group homomorphism because of our multiplicatively closed assumption and also $\pi\circ\iota=id_{G_{1}/H_{1}}$. Hence the sequence is right split. Thus $M\cong(H_{1}\times H_{2})\rtimes G_{1}/H_{1}$.
Now assume that the statement is true for $k=m$, where $2\leq m\leq l-1$, that is,
${\rm Gal}(E_{1}E_{2}\cdots E_{m}/K)\cong(H_{1}\times H_{2}\times\dots\times H_{m}),$
$\text{and},{\rm Gal}(E_{1}E_{2}\cdots E_{m}/\mathbb{Q})\cong(H_{1}\times H_{2}\times\dots\times H_{m})\rtimes G_{1}/H_{1},$
we will prove the statement for $k=m+1$.
Let $F_{1}=E_{1}\cdots E_{m}$ and $F_{2}=E_{m+1}$. Hence
$\begin{split}{\rm Gal}(F_{1}/\mathbb{Q})&={\rm Gal}(E_{1}E_{2}\cdots E_{m}/\mathbb{Q})\cong(H_{1}\times H_{2}\times\dots\times H_{m})\rtimes G_{1}/H_{1},\\ {\rm Gal}(F_{2}/\mathbb{Q})&={\rm Gal}(E_{m+1}/\mathbb{Q})\cong G_{m+1},\\ {\rm Gal}(F_{1}/K)&={\rm Gal}(E_{1}E_{2}\cdots E_{m}/K)\cong(H_{1}\times H_{2}\times\dots\times H_{m}),\\ {\rm Gal}(F_{2}/K)&={\rm Gal}(E_{m+1}/K)\cong H_{m+1}\ \text{since}F_{1}\cap F_{2}=E_{1}\cdots E_{m}\cap E_{m+1}=K.\end{split}$
Hence $F_{1}$ and $F_{2}$ satisfy conditions of base case and thus we have
$\begin{split}{\rm Gal}(E_{1}E_{2}\cdots E_{m+1}/K)&={\rm Gal}(F_{1}F_{2}/K)\cong{\rm Gal}(F_{1}/K)\times{\rm Gal}(F_{2}/K)\\ &\cong(H_{1}\times H_{2}\times\dots\times H_{m})\times H_{m+1}=H_{1}\times H_{2}\times\dots\times H_{m+1}\end{split}$
and,
$\begin{split}{\rm Gal}(E_{1}E_{2}\cdots E_{m+1}/\mathbb{Q})&={\rm Gal}(F_{1}F_{2}/\mathbb{Q})\cong{\rm Gal}(F_{1}F_{2}/K)\rtimes G_{1}/H_{1}\\ &\cong(H_{1}\times H_{2}\times\dots\times H_{m+1})\rtimes G_{1}/H_{1}.\end{split}$
Remark 2.4.1.
  1. (1)
    If $[G:H]$ is finite, then a set of representatives for $H$-cosets in $G$ satisfies the condition of closed under multiplication if and only if it is a subgroup of $G$.
    In particular, A set of representatives $\{1,x,x^{2},...,x^{n-1}\}$ for $H$-cosets in $G$ which forms a cyclic subgroup of $G$, satisfies the condition. We get a criterion for existence of such a set of representatives in Proposition 2.2.
  2. (2)
    The second condition assumed in the theorem is equivalent to exact sequences
    (2.4) $1\rightarrow H_{i}\rightarrow G_{i}\rightarrow G_{i}/H_{i}\rightarrow 1$
    being right split, that is $G_{i}\cong H_{i}\rtimes G_{i}/H_{i}$ for some semidirect product group law.
Remark 2.4.2.
Let $d_{i}$ be distinct primes for $i=1,2$ and $d_{3}=d_{1}d_{2}$. Let $K=\mathbb{Q}$ and consider quadratic extensions $E_{i}=\mathbb{Q}\sqrt{d_{i}}$ for $1\leq i\leq 3$.
Clearly $E_{i}\cap E_{j}=K$ for all $i\neq j$, but we have $E_{i}E_{j}\cap E_{k}=E_{k}\neq K$ where $i,j,k$ are a permutation of $1,2,3$. So the assumed condition in the above theorem is not always satisfied and hence it is important.
Remark 2.4.3.
In the above theorem, we could have assumed a symmetric but stronger condition
$E_{1}...E_{i-1}E_{i+1}...E_{n}\cap E_{i}=K$ for all $1\leq i\leq n$ which implies the condition that we have assumed $E_{1}...E_{k}\cap E_{k+1}=K$ for all $1\leq k\leq n-1$ since $E_{1}...E_{k}\cap E_{k+1}\subset E_{1}...E_{k}E_{k+2}...E_{n}\cap E_{k+1}=K$. But it would have made our theorem weaker. (Note: Condition for $i=1$ is not required, it is written for symmetry).
Remark 2.4.4.
  1. (1)
    Under the conditions of above theorem, we have $(H_{1}\times H_{2}\times\dots\times H_{l})\rtimes G_{1}/H_{1}\xhookrightarrow{}G_{1}\times G_{2}\times\dots\times G_{l}$ from [2, Theorem 1.1].
  2. (2)
    If all $G_{i}$s are abelian in above theorem that is each $E_{i}$ is an abelian extension of $\mathbb{Q}$, then the compositum is an abelian extension over $\mathbb{Q}$ and the semidirect product that we defined is in fact the direct product.
  3. (3)
    However to the best of our knowledge, one doesn’t know in general whether $(H_{1}\times H_{2}\times\dots\times H_{l})\times G_{1}/H_{1}$ is realizable as a Galois group over $\mathbb{Q}$.
We now describe an independently interesting consequence of the Galois Correspondence (see [2, Theorem 2.1] and [2, Theorem 2.6].)
Proposition 2.5.
Let finite groups $H_{1},H_{2},...,H_{n}$ be Galois groups over $K$ of extensions $E_{1},E_{2},...,E_{n}$ respectively which are contained in $\bar{\mathbb{Q}}$. Fix $k\leq n-1$ and suppose $E_{i_{1}}\cdots E_{i_{j}}\cap E_{i_{j+1}}=K$ for all $1\leq j\leq k-1$ where $i_{l}$ are distinct elements from $1$ to $k$. Then the following statements are equivalent.
  1. (1)
    $E_{1}\cdots E_{k}\cap E_{k+1}=K$.
  2. (2)
    $E_{i_{1}}\cdots E_{i_{k}}\cap E_{i_{k+1}}=K$ where $i_{l}$ are distinct elements from $1$ to $k+1$.
  3. (3)
    $E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{k+1}\cap E_{1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}=E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1}$ for any $1\leq i<j\leq k+1$.
Proof.
Because of given conditions and induction we get, ${\rm Gal}(E_{i_{1}}\cdots E_{i_{j}}/K)\cong H_{i_{1}}\times...\times H_{i_{j}}$ for all $j\leq k$ where $i_{l}$ are distinct elements from $1$ to $k$ and ${\rm Gal}(E_{i_{1}}\cdots E_{i_{j}}/E_{i_{1}}\cdots E_{i_{m-1}}E_{i_{m+1}}\cdots E_{i_{j}})\cong H_{i_{m}}$ for all $j\leq k$ and $1\leq m\leq j$ where $i_{l}$ are distinct elements from $1$ to $k$.
Equivalence of (1) and (2): For $i_{l}$ distinct elements from $1$ to $k+1$,
$E_{1}...E_{k}\cap E_{k+1}=K$.
$\iff$ ${\rm Gal}(E_{1}\cdots E_{k+1}/K)\cong H_{1}\times\dots\times H_{k+1}$.
$\iff$ $E_{i_{1}}\cdots E_{i_{k}}\cap E_{i_{k+1}}=K$.
Equivalence of (1) and (3): For any $1\leq i<j\leq k+1$, let
$\begin{split}D&=E_{1}\cdots E_{k+1},\\ D_{i}&=E_{1}\cdots E_{i-1}E_{i+1}\cdots E_{k+1},\\ D_{j}&=E_{1}\cdots E_{j-1}E_{j+1}\cdots E_{k+1},\\ D_{ij}&=E_{1}...E_{i-1}E_{i+1}...E_{j-1}E_{j+1}...E_{k+1}.\end{split}$
Now since ${\rm Gal}(D_{i}/D_{ij})\cong H_{j}$ and ${\rm Gal}(D_{j}/D_{ij})\cong H_{i}$, we have by [2, Theorem 1.1],
${\rm Gal}(D/D_{ij})\xhookrightarrow{}H_{i}\times H_{j}.$
We also have
$\begin{split}{\rm Gal}(D/K)/{\rm Gal}(D/D_{ij})&\cong{\rm Gal}(D_{ij}/K)\\ &\cong H_{1}\times\dots\times H_{i-1}\times H_{i+1}\times\dots\times H_{j-1}\times H_{j+1}\times\dots\times H_{k+1}.\end{split}$
Hence,
$\begin{split}&{\rm Gal}(D/K)\cong H_{1}\times\dots\times H_{k+1}\\ \implies&{\rm Gal}(D/D_{ij})\cong H_{i}\times H_{j}\\ \implies&D_{i}\cap D_{j}=D_{ij}.\end{split}$
Conversely,
$\begin{split}&D_{i}\cap D_{j}=D_{ij}\\ \implies&D_{i}\cap E_{i}\subset D_{i}\cap D_{j}=D_{ij}\\ \implies&D_{i}\cap E_{i}\subset D_{ij}\cap E_{i}=K\\ \implies&D_{i}\cap E_{i}=K\\ \implies&{\rm Gal}(D/K)\cong H_{1}\times\cdots\times H_{k+1}.\end{split}$
2.3. Application to cases in work of Arias-de-Reyna & König
Theorem 2.6.
Let $n\in\mathbb{N}$. For $q$ suppose the following hold.
  1. (1)
    $q\equiv 1$ modmod $n$ and (n,(q1)/n)=1(n,(q-1)/n)=1.
  2. (2)
    The group ${\rm GL}_{2}(\mathbb{F}_{q})$ is realizable as Galois group over $\mathbb{Q}$ of extensions $E_{1},E_{2}$ which are contained in $\bar{\mathbb{Q}}$ such that $E_{1}\cap E_{2}=K$ is a degree $n$ abelian extension of $\mathbb{Q}$.
Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{q})/H$ (with semidirect product group law as in Theorem 2.4) is realizable as Galois group over $\mathbb{Q}$ where $H$ is the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
Proof.
We have for $l=2$, $E_{1},E_{2}$ satisfying conditions of Theorem 2.4 with $G_{1}=G_{2}={\rm GL}_{2}(\mathbb{F}_{q})$, and $G_{1}:H_{1}]=n$ and $G_{1}/H_{1}$ is abelian group and $q\equiv 1$ modmod $n$. Therefore $H_{1}=H_{2}=H$. Since $q\equiv 1$ modmod $n$ and (n,(q1)/n)=1(n,(q-1)/n)=1, from Corollary 2.2.1, we get the required set of representatives for $H$-cosets in ${\rm GL}_{2}(\mathbb{F}_{q})$.∎
Remark 2.6.1.
Because of the conditions on $p$ and $n$ in above theorem, we actually get that $K$ is a cyclic extension (not just abelian).
If extensions $E_{1},E_{2}$ of $\mathbb{Q}$ contained in $\bar{\mathbb{Q}}$, are linearly disjoint over an extension $K$ of $\mathbb{Q}$, then $E_{1}\cap E_{2}=K$. We get from [1, Remark 4.4] and [1, Corollary 4.3], that there are two Galois extensions $L_{1},L_{2}$ over $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$ which are linearly disjoint over $\mathbb{Q}(\zeta_{p})$.
Corollary 2.6.1.
For a prime $p\geq 5$, $(SL_{2}(\mathbb{F}_{p})\times SL_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/(p-1)\mathbb{Z}$ (with semidirect product group law as in above theorem) is realizable as Galois group over $\mathbb{Q}$.
Proof.
From [1, Remark 4.4] and [1, Corollary 4.3], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with n=p1n=p-1, $K=\mathbb{Q}(\zeta_{p})$ and ${\rm GL}_{2}(\mathbb{F}_{p})/H$ is cyclic group of order p1p-1. Moreover from Corollary 2.1.2 (2), we have $H=SL_{2}(\mathbb{F}_{p})$. ∎
From previous corollary and proof of Theorem 2.4 we have the following.
Corollary 2.6.2.
For a prime $p\geq 5$, $SL_{2}(\mathbb{F}_{p})$ and $SL_{2}(\mathbb{F}_{p})\times SL_{2}(\mathbb{F}_{p})$ are realizable as Galois groups over $\mathbb{Q}(\zeta_{p})$.
Remark 2.6.2.
It is proved in [1, Theorem 1.1] that for a prime $p\geq 5$, there are infinitely many locally cyclic Galois extensions of $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$, which are pairwise linearly disjoint over $\mathbb{Q}(\sqrt{p^{*}})$ where $p^{*}=(-1)^{(p-1)/2}p$.
Corollary 2.6.3.
For a prime $p\geq 5$ with $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ (hence normal) subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{p})/H$ (with semidirect product group law as in above theorem) is realizable as Galois group over $\mathbb{Q}$.
Proof.
From [1, Theorem 1.1], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with $n=2$. Since $p\equiv 1$ modmod 22, the conditions $p\equiv 3$ modmod 44 and (2,(p1)/2)=1(2,(p-1)/2)=1 are equivalent. Hence we are done. ∎
Remark 2.6.3.
If $p\equiv 1$ modmod 44 then (2,(p1)/2)=2(2,(p-1)/2)=2. Hence by Corollary 2.2.1 there is no $x\in{\rm GL}_{2}(\mathbb{F}_{p})$ such that $x\in H$ and $order(x)=2$.
Now for $p\equiv 3$ modmod 44 we have $p^{*}=-p$. From previous corollary and proof of Theorem 2.4 we have the following.
Corollary 2.6.4.
For a prime $p\geq 5$ with $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $H$ and $H\times H$ are realizable as Galois groups over $\mathbb{Q}\sqrt{-p}$.
Proposition 2.7.
For a prime $p\geq 5$, $(PSL_{2}(\mathbb{F}_{p})\times PSL_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/2\mathbb{Z}$ (with semidirect product group law as in Theorem 2.4) is realizable as Galois group over $\mathbb{Q}$.
Proof.
Consider $H=\{x\in{\rm GL}_{2}(\mathbb{F}_{p})\ |\ det(x)$ is a square in $\mathbb{F}_{p}^{\times}\}$, which is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ by Corollary 2.1.1. It can be seen that $H=Z({\rm GL}_{2}(\mathbb{F}_{p}))SL_{2}(\mathbb{F}_{p})$. Hence it follows that $PH\cong PSL_{2}(\mathbb{F}_{p})$. Since $H$ is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$ containing $Z({\rm GL}_{2}(\mathbb{F}_{p}))$, we have that $PH\cong PSL_{2}(\mathbb{F}_{p})$ is the unique index-$2$ subgroup of ${\rm PGL}_{2}(\mathbb{F}_{p})$.
We have $p\geq 5$. Let $r\neq p-1$ be a non-square element in $\mathbb{F}_{p}$. Consider $x=\begin{pmatrix}1&r+1\\ -1&-1\end{pmatrix}\not\in H$ with $x^{2}=1\ (Z({\rm GL}_{2}(\mathbb{F}_{p})))$.
From [1], we have that [1, Theorem 1.1] is also true for ${\rm PGL}_{2}(\mathbb{F}_{p})$. Thus we get $E_{1},E_{2}$ satisfying conditions of Theorem 2.4 for $l=2$ with $G_{1}=G_{2}={\rm PGL}_{2}(\mathbb{F}_{p})$, and $G_{1}:H_{1}]=2$. Therefore $H_{1}=H_{2}=PSL_{2}(\mathbb{F}_{p})$. We observe that $\{1,xZ({\rm GL}_{2}(\mathbb{F}_{p}))\}$ is required set of representatives of $PSL_{2}(\mathbb{F}_{p})$-coset in ${\rm PGL}_{2}(\mathbb{F}_{p})$. ∎
Corollary 2.7.1.
For a prime $p\geq 5$, $PSL_{2}(\mathbb{F}_{p})$ and $PSL_{2}(\mathbb{F}_{p})\times PSL_{2}(\mathbb{F}_{p})$ are realizable as Galois groups over $\mathbb{Q}\sqrt{p^{\ast}}$.
3. Galois Representations, Right Splitting and Inverse Galois Problem
In this section, by using the algebraic operations induction, direct sums and tensor products on Galois representations and right splitting of some exact sequences of groups, we establish occurrence of some groups as Galois groups over $\mathbb{Q}$}.
3.1. Induced Galois representations and Galois groups
Definition 3.1.
Let $G={\rm Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$. Let $E$ be a number field and $\Lambda$ be a prime in its ring of integers $\mathcal{O}$. Let $E_{\Lambda}$ be the completion of $E$ with respect to $\Lambda$ and $\mathcal{O}_{\Lambda}$ be its ring of integers. Let $\mathbb{F}_{\Lambda}=\mathcal{O}_{\Lambda}/\Lambda\mathcal{O}_{\Lambda}$.
We call $(\pi,V)$ a Galois representation if $V$ is a finite dimensional vector space over $E_{\Lambda}$ and $\pi:G\rightarrow{\rm GL}(V)$ is a continuous homomorphism.
Remark 3.1.1.
Let $dim_{E_{\Lambda}}(V)=m$. Now from [3, Proposition 9.3.5], there is a basis of $V$ such that $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$ for all $g\in G$. Let us consider the maps $\pi^{\prime}:G\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$ obtained from $\pi$ through the quotients maps $\mathcal{O}_{\Lambda}\rightarrow\mathbb{F}_{\Lambda}$ and ${\rm GL}_{m}(\mathbb{F}_{\Lambda})\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$.
Definition 3.2.
[…]
We have a similar definition for $\pi^{\prime}$ in place of $\tilde{\pi}$.
Let $K$ be a finite extension of $\mathbb{Q}$ contained in $\bar{\mathbb{Q}}$. Then $\bar{K}=\bar{\mathbb{Q}}$. Let $H={\rm Gal}(\bar{K}/K)\subset G$. Let $(\pi,W)$ be a Galois representation with $dim_{E_{\Lambda}}(W)=m$.
Let $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$. Then $\sigma^{\prime}:H\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\sigma}:H\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$ and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)$. Also, $\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]$.
The following discussion is valid for a general context of a finite dimensional representation of a finite index subgroup $H$ of a group $G$. In this manuscript, we only consider the case of ${\rm Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$ and ${\rm Gal}(\bar{K}/K)$.
Let us consider the induced representation $\rho=Ind_{H}^{G}(\sigma)$ on the induced space over $E_{\Lambda}$, defined by
$V=\{f:G\rightarrow W\ |\ f(hg)=\sigma(h)f(g)\ for\ all\ h\in H,g\in G\}$
and $\rho:G\rightarrow{\rm GL}(V)$ is given by $\rho(g)f(x)=f(xg),\text{for all}\ g,x\in G,f\in V$. Thus $(\rho,V)$ is a Galois representation.
[…]
$\phi_{s,w}(g)=\begin{cases}\sigma(gs^{-1})w,&\text{if}\ gs^{-1}\in H\\ 0&\text{otherwise.}\end{cases}$
and $dim(V)=dim(W)[G:H]$.
Let $H$ be a normal subgroup of $G$ (i.e., $K$ is Galois over $\mathbb{Q}$}) such that $[G:H]=n$. Since $dim(W)=m$, we have $dim(V)=mn$. Without loss of generality, let $s_{1}=1$. We label the basis of $V$ as
$\begin{split}f_{1}=\phi_{1,w_{1}},\ f_{2}=\phi_{1,w_{2}},\ \dots,\ f_{m}=\phi_{1,w_{m}},\\ f_{m+1}=\phi_{s_{2},w_{1}},\ \dots,\ f_{2m}=\phi_{s_{2},w_{m}},\dots,\\ f_{(n-1)m+1}=\phi_{s_{n},w_{1}},\ \dots,\ f_{nm}=\phi_{s_{n},w_{m}}.\end{split}$
We write down the matrix of $\rho(g)$ with respect to the above basis. Observe that
$\rho(g)f_{i}(x)=f_{i}(xg)=\sum\limits_{j=1}^{mn}a_{ji}f_{j}(x)$
By taking $x=s_{k}$, we get that
$\rho(g)=\begin{bmatrix}(f_{1}(g))&(f_{2}(g))&\dots&(f_{nm}(g))\\ (f_{1}(s_{2}g))&(f_{2}(s_{2}g))&\dots&(f_{nm}(s_{2}g))\\ \dots&\dots&\dots&\dots\\ (f_{1}(s_{n}g))&(f_{2}(s_{n}g))&\dots&(f_{nm}(s_{n}g))\end{bmatrix}_{nm\times nm}$
where $(f_{i}(s_{k}g))$ are treated as $m\times 1$ column matrices (matrices with respect to given basis $\{w_{j}:1\leq j\leq m\}$ of $W$).
The $nm\times nm$ matrix for $\rho(hs_{i})$: The $p,q$-th $m\times m$ block where $1\leq p,q\leq n$ and $q$ is such that $s_{p}s_{i}\in s_{q}H$ is $\pi(s_{p}hs_{i}s_{q}^{-1})$, since $\phi_{s_{q},w}(s_{p}hs_{i})=\pi(s_{p}hs_{i}s_{q}^{-1})w$, since $s_{p}hs_{i}\in s_{q}H$.
Since $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$, we conclude that $\rho(G)\subset{\rm GL}_{nm}(\mathcal{O}_{\lambda})$ and $\rho^{\prime}:G\rightarrow{\rm GL}_{nm}(\mathbb{F}_{\Lambda})$ and $\tilde{\rho}:G\rightarrow{\rm PGL}_{nm}(\mathbb{F}_{\Lambda})$ can be defined.
Remark 3.2.1.
All the following results for the pair ($\rho$, $\pi$) in this subsection also hold for pairs ($\rho^{\prime}$, $\pi^{\prime}$) as well as ($\tilde{\rho}$, $\tilde{\pi}$).
Lemma 3.3.
Images $\rho(H)$ and $\pi(H)$ are isomorphic.
Proof.
If $g=h\in H$, then
$\rho(h)=\begin{bmatrix}\pi(h)&0&\dots&\\ 0&\pi(s_{2}hs_{2}^{-1})&&\\ \vdots&&\ddots&\\ &&&\pi(s_{n}hs_{n}^{-1})\\ \end{bmatrix}_{nm\times nm}.$
Define a map from $\rho(H)$ to $\pi(H)$ by first block projection sending $\rho(h)$ to $\pi(h)$. ∎
Lemma 3.4.
We get an exact sequence
$1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1.$
Proof.
If $g=hs_{j}\in Hs_{j}$, $(f_{(j-1)m+1}(g))=\pi(h)w_{1},(f_{(j-1)m+2}(g))=\pi(h)w_{2},\dots,(f_{jm}(g))=\pi(h)w_{m}$ and for other $k$, $(f_{k}(g))=0$. We have
$\rho(G)=\rho(H)\bigsqcup\rho(Hs_{2})\bigsqcup\dots\bigsqcup\rho(Hs_{n})=\rho(H)\bigsqcup\rho(H)\rho(s_{2})\bigsqcup\dots\bigsqcup\rho(H)\rho(s_{n})$
hence $\rho(G):\rho(H)]=n$.
Now, since $H$ is normal in $G$, its left and right cosets coincide. Hence we can define a map $\gamma:\rho(G)\rightarrow G/H$ with $\gamma(\rho(hs_{k}))=s_{k}H$ for any $h,s_{k}$. Now $\gamma$ is well defined surjective homomorphism because $\rho(Hs_{i})$ for $1\leq i\leq n$ are disjoint. The kernel of $\gamma$ is precisely $\rho(H)$. ∎
Remark 3.4.1.
We do not necessarily get a similar exact sequence involving $\pi(H)$, $\pi(G)$ and $G/H$ a similar well defined surjective map from $\pi(G)$ to $G/H$ may not exist since $\pi(Hs_{i})$ need not be disjoint.
Example 3.5.
Let $G/H$ be cyclic with representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$. If $g=hs^{i}\in H$ for $0\leq i\leq n-1$, the matrix $\rho(hs^{i})$ is given by
$\begin{bmatrix}0&\dots&&0&\pi(h)&0&\dots&\\ \vdots&&&&0&\pi(shs^{-1})&0&\dots\\ &\ddots&&&\vdots&&\ddots&\\ &&&&&&&\pi(s^{n-1-i}hs^{-(n-1-i)})\\ \pi(s^{n-i}hs^{i})&&&&&&&0\\ 0&\pi(s^{n-(i-1)}hs^{(i-1)})&&&&&&\vdots\\ \vdots&0&\ddots&&&&&\\ &&0&\pi(s^{n-1}hs)&&&&\\ \end{bmatrix}$.
In particular,
$\rho(s)=\begin{bmatrix}0&I_{2}&0&\dots&\\ 0&0&I_{2}&0&\dots\\ \vdots&\vdots&&\ddots&\\ &&&&I_{2}\\ \pi(s^{n})&&&&0\\ \end{bmatrix}$.
By computing and comparing the matrices $\rho(s_{i}),\rho(s_{j})$, $\rho(s_{k})$ and $\rho(s_{i})\rho(s_{j})$, we get the following.
Lemma 3.6.
Suppose we have a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$, then $\rho(s_{i})\rho(s_{j})=\rho(s_{k})\iff(\pi(s_{i})\pi(s_{j})=\pi(s_{k})$ and $s_{i}s_{j}\in s_{k}H$).
Remark 3.6.1.
$\pi(s_{i})$ need not be distinct even though $\rho(s_{i})$ are distinct. In fact even if all $\pi(s_{i})$ are same, $\rho(s_{i})$ will be distinct.
Corollary 3.6.1.
Let $G/H$ be cyclic with representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$. Then $\rho(s)^{n}=1$ if and only if $\pi(s)^{n}=1$.
Theorem 3.7.
Let $G$ and $H$ be as above, then the exact sequence
(3.1) $1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1$
is right split (that is $\rho(G)\cong\rho(H)\rtimes G/H$ for some semidirect product group law) if and only if there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$. (In fact it is a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H$)
Proof.
Let $\{r_{1},r_{2},...,r_{n}\}$ be a set of representatives of $H$-cosets in $G$. If the exact sequence is right split, let $\iota$ be splitting with $\gamma\circ\iota=id_{G/H}$. Hence $\gamma\circ\iota(r_{i}H)=r_{i}H$ for each $i$. Thus for each $i$, $\iota(r_{i}H)=\rho(h_{i}r_{i})$ for some $h_{i}\in H$. Let $s_{i}=h_{i}r_{i}$ for each $i$. Hence $\{s_{1},s_{2},...,s_{n}\}$ also forms a set of representatives for $H$-cosets in $G$ with $s_{i}H=r_{i}H$ and $\iota(s_{i}H)=\rho(s_{i})$ for each $i$. Since $\iota$ is a homomorphism, $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$.
Conversely, suppose there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$ and $\{\rho(s_{i})\}_{i}$ forming a multiplicatively closed subset of $\rho(G)$. Hence $\{\rho(s_{i})\}_{i}$ forms a subgroup of $\rho(G)$. $\{\rho(s_{i})\}_{i}$ is a group with the same group structure as $\{s_{i}H\}_{i}=G/H$. Then we can define $\iota(s_{i}H)=\rho(s_{i})$. Hence $\iota$ is a homomorphism. Also, $\gamma(\iota(s_{i}H))=\gamma(\rho(s_{i}))=s_{i}H$ hence $\gamma\circ\iota=id_{G/H}$. Hence the exact sequence is right split. ∎
Remark 3.7.1.
We can similarly prove the statements in Remark 2.4.1 in previous section.
Corollary 3.7.1.
If there exists a set of coset representatives of $H$ in $G$ which forms a multiplicatively closed set, then the exact sequence 3.1 is right split.
Corollary 3.7.2.
Let $K$ be a cyclic extension of $\mathbb{Q}$ and let $G$ and $H$ be as above. Then the exact sequence 3.1 is right split if and only if there exists a set of representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$ with $\rho(s)^{n}=1$. Also, if these statements are true then $order(\rho(s))=n$ and $s^{n}\in H$.
Proof.
From the above theorem, we get coset representatives of $H$, $\{s_{i}\}_{i}$ with $\{\rho(s_{i})\}_{i}$ forming a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H=<sH>$ for some $s\in G$ with $s_{1}\in H$. Hence without loss of generality, let $s_{i}H=s^{i-1}H$ for all $i$. Now let $r_{i}=s_{2}^{i-1}$. Hence $r_{i}H=s_{i}H$. Now since $s_{2}^{n}H=(s_{2}H)^{n}=H$, we have $\tilde{\rho}(s_{2})^{n}=1$. Hence we are done. Other assertions are clear.∎
Applying above discussion to $\tilde{\rho}$ and $\tilde{\pi}$ we have the following.
Theorem 3.8.
Let $H$ be a finite index normal subgroup of $G=Gal(\bar{\mathbb{Q}}/\mathbb{Q})$. Suppose there exists a set of coset representatives of $H$ in $G$ which form a multiplicatively closed subset of $G$. If a finite group $M$ is realizable as a Galois group over $\mathbb{Q}$ through a Galois Representation $(\pi,W)$ such that $M\cong\tilde{\pi}(G)=\tilde{\pi}(H)$, then $M\rtimes G/H$ is realisable as a Galois group over $\mathbb{Q}$ (for semidirect product group law as in Theorem 3.7).
Observation : $G=H\iff ker(\pi)\subset H\ and\ \pi(G)=\pi(H).$
We can generalize Lemma  3.4 and Theorem 3.7. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
Theorem 3.9.
  1. (1)
    We get an exact sequence
    (3.2) $1\rightarrow\rho(H^{\prime}\cap H)\rightarrow\rho(H^{\prime})\rightarrow H^{\prime}/(H^{\prime}\cap H)\rightarrow 1.$
  2. (2)
    The exact sequence 3.2 is right split (that is $\rho(H^{\prime})\cong\rho(H^{\prime}\cap H)\rtimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law) if and only if there exists a set of representatives of $(H^{\prime}\cap H)$-cosets in $H^{\prime}$, $\{s_{1},s_{2},\dots,s_{l}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(H^{\prime})$.
  3. (3)
    $H\subset H^{\prime}$ if and only if $\rho(H)\subset\rho(H^{\prime})$ and $ker(\pi)\cap H\subset H^{\prime}$
  4. (4)
    Let $H\subset H^{\prime}$. Then the exact sequence 3.2 is 1ρ(H)ρ(H’) H’/H 1. If there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$, then this exact sequence is right split. That is $\rho(H^{\prime})\cong\rho(H)\rtimes H^{\prime}/H$ for some semidirect product group law.
3.2. Galois representations for newforms and application to the cases in work of Zywina
Let $p\geq 5$. Fix non-CM newform $f$ of weight $k>1$ on $\Gamma_{1}(N)$. Let $l$ be rational prime as in [8]. Let $\pi$ be the Deligne representation associated with $f$ with properties as in [8, Section 1.1]. Let $W=E_{\Lambda}^{2}$ and $\pi:G\rightarrow{\rm GL}(W)$ with $\pi(G)\subset{\rm GL}_{2}(\mathcal{O}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$.
Let $p=l$ and assume conditions of [8, Section 1.2]. Then ${\rm Image}(\tilde{\pi})=\tilde{\pi}(G)={\rm PSL}_{2}(\mathbb{F}_{p})$. The representations $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$ and $\tilde{\sigma}:H\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$ are defined, and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)\subset\tilde{\pi}(G)$.
Lemma 3.10.
Let $p\geq 5$ and $H$ be normal in $G$ (i.e. K is Galois over Q) and $G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Then $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
Proof.
Now, ${\rm PSL}_{2}(\mathbb{F}_{p}):\tilde{\pi}(H)]=[\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence $\tilde{\pi}(H)$ is not trivial. Since $H$ is normal in $G$, $\tilde{\pi}(H)$ is normal in $\tilde{\pi}(G)$. Now since ${\rm PSL}_{2}(\mathbb{F}_{p})$ is simple we have that ${\rm PSL}_{2}(\mathbb{F}_{p})=\tilde{\pi}(H)$ otherwise $\tilde{\pi}(H)$ will become non-trivial normal subgroup in ${\rm PSL}_{2}(\mathbb{F}_{p})$.∎
Remark 3.10.1.
If $n\geq|{\rm PSL}_{2}(\mathbb{F}_{p})|$ then either $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$ or $\tilde{\pi}(H)$ is trivial. If latter case happens, then $\tilde{\rho}(H)$ is also trivial. Hence $\tilde{\rho}(G)\cong G/H$ in that case.
Proposition 3.11.
  1. (1)
    For $p\geq 5$, ${\rm PGL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ for some semidirect product group law and the semidirect product is not direct product.
  2. (2)
    Let $p\geq 5$. For any semidirect product group law such that ${\rm PGL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$, the semidirect product is not direct product. Furthermore, the automorphism of ${\rm PSL}_{2}(\mathbb{F}_{p})$, given by conjugation by image of generator of $\mathbb{Z}/2\mathbb{Z}$ in ${\rm PGL}_{2}(\mathbb{F}_{p})$, is not inner.
Theorem 3.12.
  1. (1)
    For $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7).
  2. (2)
    This semidirect product in part (1) is direct $\iff$ $\tilde{\pi}(s)=I$.
  3. (3)
    Automorphism $\phi_{\tilde{\rho}(s)}$ of $\tilde{\rho}(H)$, by conjugation by $\tilde{\rho}(s)$, is inner.
  4. (4)
    The group obtained here is not isomorphic to ${\rm PGL}_{2}(\mathbb{F}_{p})$.
Proof.
(1) Let $[G:H]=2$. Then $H$ is normal in $G$ ($K$ is Galois). Also $2<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence by Lemma 3.10, $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
We could have chosen $K=\mathbb{Q}(i)$ and $s\in G$ as complex conjugation so that $s^{2}=1$ and $\{1,s\}$ become representatives of right cosets of $H$ in $G$. Then by Theorem 3.8 we are done.
(2) Now, $\tilde{\pi}(G)=\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
Above semidirect product is direct.
$\iff$ $\tilde{\rho}(s)$= $\begin{bmatrix}0&I_{2}\\ I_{2}&0\end{bmatrix}$ commutes with every $\tilde{\rho}(h)=$ $\begin{bmatrix}\tilde{\pi}(h)&0\\ 0&\tilde{\pi}(shs^{-1})\end{bmatrix}$.
$\iff$ $\tilde{\pi}(s)$ commutes with every $\tilde{\pi}(h)$.
$\iff$ $\tilde{\pi}(s)\in Z({\rm PSL}_{2}(\mathbb{F}_{p}))$.
$\iff$ $\tilde{\pi}(s)=I$.
Since $Z({\rm PSL}_{2}(\mathbb{F}_{p}))$ is trivial as ${\rm PSL}_{2}(\mathbb{F}_{p})$ is non-abelian and simple.
(3) Now $\tilde{\pi}(G)=\tilde{\pi}(H)$. Hence $\tilde{\pi}(s)=\tilde{\pi}(h^{\prime})$ for some $h^{\prime}\in H$.
So, $\phi_{\tilde{\rho}(s)}(\tilde{\rho}(h))=\tilde{\rho}(s)\tilde{\rho}(h)\tilde{\rho}(s)^{-1}=$ $\begin{bmatrix}\tilde{\pi}(shs^{-1})&0\\ 0&\tilde{\pi}(h)\end{bmatrix}$= $\tilde{\rho}(h^{\prime})\tilde{\rho}(h)\tilde{\rho}(h^{\prime})^{-1}$.
(4) This follows from (3) and Proposition 3.11.∎
We have results similar to (1), (2) and (3) of Theorem 3.12 for certain simple groups of the form ${\rm PSL}_{2}(\mathbb{F}_{q})$.
Theorem 3.13.
The group ${\rm PSL}_{2}(\mathbb{F}_{q})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7) for following $q$.
  1. (1)
    $q=p$ for $p\geq 5$ (From [8, Section 1.2]).
  2. (2)
    $q=p^{2}$ for $p\equiv\pm 2\ mod\ 5,\ p\geq 7$ (From [5, Corollary 3.6]).
  3. (3)
    $q=p^{2}$ for $p\equiv\pm 3\ mod\ 8,\ p\geq 5$ (From [5, Corollary 3.8]).
  4. (4)
    $q=p^{3}$ for odd prime $p\equiv\pm 2,\pm 3,\pm 4,\pm 6\ mod\ 13$ (From [8, Section 1.3]).
  5. (5)
    $q=5^{3},3^{5},3^{4}$ (From [4, Section 2.2, 2.5, 3]).
Remark $3.1$.
One has even more general conditions for $q=p^{2}$ (See [6, Theorem 3.1]) and $q=p^{4}$ (See [6, 3.3]).
3.3. Direct sum / tensor product of representations and Galois groups
Lemma 3.14.
If for each $1\leq i\leq n$, finite group $H_{i}$ is realizable as a Galois group over $\mathbb{Q}$ through Galois representation $(\pi_{i},V_{i})$ for $m_{i}$-dimensional vector spaces $V_{i}$ over $E_{\Lambda}$, then $\{(\tilde{\pi}_{1}(g),...,\tilde{\pi}_{n}(g))\ |\ g\in G\}\subset H_{1}\times H_{2}\times\dots\times H_{n}$ is realisable as Galois group over $\mathbb{Q}$.
Proof.
[…]
We consider the tensor product Galois representation $\Pi=\bigotimes\limits_{1\leq i\leq n}\pi_{i}$ of $G$ on $\bigotimes\limits_{1\leq i\leq n}V_{i}$.
Proposition 3.15.
The groups realized as Galois groups over $\mathbb{Q}$ through Galois representations $\pi$ and $\Pi$ are isomorphic, i.e.,
[…]
Proof.
Let $V_{1}$ and $V_{2}$ be finite dimensional vector spaces over a general field $\mathbb{F}$} with dimensions $m_{1}$ and $m_{2}$ respectively. Given $T_{i}\in{\rm GL}(V_{i})$ for each $i=1,2$, we consider the natural $\mathbb{F}$}-linear map $(T_{1}\otimes T_{2}):V_{1}\otimes V_{2}\rightarrow V_{1}\otimes V_{2}$ that is also invertible. Fix bases of $V_{1}$ and $V_{2}$ and the corresponding basis of $V_{1}\otimes V_{2}$. Let $A_{1},A_{2},A$ be the matrices representing $T_{1},T_{2},T_{1}\otimes T_{2}$, respectively with respect to these bases. Then
$A=A_{1}\otimes A_{2}=\begin{bmatrix}a_{11}A_{2}&\dots&a_{1m_{1}}A_{2}\\ \vdots&\ddots&\vdots\\ a_{m_{1}1}A_{2}&\dots&a_{m_{1}m_{1}}A_{2}\end{bmatrix}_{m_{1}m_{2}},$
where $A_{1}=\begin{bmatrix}a_{11}&\dots&a_{1m_{1}}\\ \vdots&\ddots&\vdots\\ a_{m_{1}1}&\dots&a_{m_{1}m_{1}}\end{bmatrix}$ and $A_{2}=\begin{bmatrix}a^{\prime}_{11}&\dots&a^{\prime}_{1m_{2}}\\ \vdots&\ddots&\vdots\\ a^{\prime}_{m_{2}1}&\dots&a^{\prime}_{m_{2}m_{2}}\end{bmatrix}$.
If the map $T_{1}\otimes T_{2}$ is given by the scalar multiplication by $\lambda\in{\mathbb{F}}^{\times}$, then it follows that $A_{i}=\mu_{i}I$ for some scalars $\mu_{1},\mu_{2}\in{\mathbb{F}}^{\times}$ such that $\lambda=\mu_{1}\mu_{2}$. Thus the map $T_{1}\otimes T_{2}$ descends to an injective group homomorphism
$\tilde{\tau}:{\rm PGL}(V_{1})\times{\rm PGL}(V_{2})\rightarrow{\rm PGL}(V_{1}\otimes V_{2}).\vskip 3.0pt plus 1.0pt minus 1.0pt$
We take the special cases ${\mathbb{F}}=E_{\Lambda}$ and then ${\mathbb{F}}=\mathbb{F}_{\Lambda}$ to complete the proof.
Let ${\mathbb{F}}=E_{\Lambda}$. Then $\Pi(g)=\pi_{1}(g)\otimes\pi_{2}(g)\otimes\dots\otimes\pi_{n}(g)$. Hence $\Pi(G)\subset{\rm GL}_{M}(\mathcal{O}_{\Lambda})$ where $M=m_{1}m_{2}\cdots m_{n}$ and $\tilde{\Pi}:G\rightarrow{\rm PGL}_{M}(\mathbb{F}_{\Lambda})$.
[…]
By induction on $n$, we get a well defined injective homomorphism from ${\rm PGL}(V_{1})\times{\rm PGL}(V_{2})\times\dots\times{\rm PGL}(V_{n})$ to ${\rm PGL}(\bigotimes\limits_{1\leq i\leq n}V_{i})$ such that ${\rm Image}(\tilde{\Pi})\cong\{(\tilde{\pi}_{1}(g),...,\tilde{\pi}_{n}(g))\ |\ g\in G\}$.∎
Let $L=\{(\pi(g),\rho(g))|g\in G\}$. Thus $\tilde{L}=\{(\tilde{\pi}(g),\tilde{\rho}(g))|g\in G\}$ is realisable as a Galois group over $\mathbb{Q}$ through the Galois representation $(\pi\oplus\rho,W\oplus V)$ where $\rho$ is induced representation as before.
Remark 3.15.1.
All the following results for the pair ($\rho$, $\pi$) also hold for pairs ($\rho^{\prime}$, $\pi^{\prime}$) as well as ($\tilde{\rho}$, $\tilde{\pi}$).
Lemma 3.16.
For $H\unlhd G$ with $[G:H]=n$, we define the projection map $\Psi:L\rightarrow\rho(G)$ by $\Psi((\pi(g),\rho(g)))=\rho(g)$ and the projection map $\Phi:L\rightarrow\pi(G)$ by $\Phi((\pi(g),\rho(g)))=\pi(g)$. Then
  1. (1)
    The map $\Psi$ is an isomorphism.
  2. (2)
    We have $ker(\Phi)\cong ker(\pi)/(ker(\pi)\cap H)$ and it can be seen as a subgroup of $G/H$.
Proof.
(1) If $\rho(g)=I$, then $g\in H$. By matrix of $\rho(g)$, we get $\pi(g)=I$ since $g\in H$. Hence $\Psi$ is also injective.
(2) Suppose $\pi(h_{i}s_{i})=\pi(h_{j}s_{j})=I_{m}$ for some $h_{i},h_{j}\in H$. Then in the matrix of $\rho(h_{i}s_{i})$, the $p,q$-th $m\times m$ block matrix, where $1\leq p,q\leq n$ and $q$ is such that $s_{p}s_{i}\in s_{q}H$, is $\pi(s_{p}s_{q}^{-1})$. Hence it is independent of $h_{i}\in H$. Let $\rho(h_{i}s_{i})=\theta_{i}$. Similarly, $\rho(h_{j}s_{j})=\theta_{j}$.
Let $s_{i}s_{j}H=s_{k}H$ for some $k$. Hence for some $h_{k}\in H$, $\pi(h_{i}s_{i})\pi(h_{j}s_{j})=\pi(h_{k}s_{k})$ and $\rho(h_{i}s_{i})\rho(h_{j}s_{j})=\rho(h_{k}s_{k})$. Hence $\pi(h_{k}s_{k})=I_{m}$ for above $h_{k}$. Let $\rho(h_{k}s_{k})=\theta_{k}$. Then $\theta_{i}\theta_{j}=\theta_{k}$. Hence $\{\theta_{i}\}_{1\leq i\leq n}$ form a group with same group law as $\{(s_{i}H)\}_{1\leq i\leq n}$. Also, for any $i$, $\rho(h_{i}s_{i})=\theta_{i}$ if and only if $\pi(h_{i}s_{i})=I_{m}$. Let $J=\{i\in\{1,2,...,n\}\ |\ ker(\pi)\cap Hs_{i}\neq\emptyset\}$. Hence $ker(\Phi)=\{(I,\theta_{i})\}_{i\in J}$. Thus we have $\Omega:ker(\Phi)\xhookrightarrow{}G/H$ sending $(I,\theta_{i})$ to $s_{i}H$.
We have the usual maps $ker(\pi)\xhookrightarrow{}G\rightarrow G/H$. Hence $ker(\pi)/(ker(\pi)\cap H)\xhookrightarrow{}G/H$. Also, $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap Hs_{i})$. Now for $i\in J$ there exists an $h_{i}\in H$ such that $\pi(h_{i}s_{i})=I_{m}$. Hence for $i\in J$, $(ker(\pi)\cap Hs_{i})=(ker(\pi)\cap H)(h_{i}s_{i})$. Hence $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap H)h_{i}s_{i}$. Thus
$ker(\pi)/(ker(\pi)\cap H)\cong\{(s_{i}H)\}_{i\in J}\cong ker(\Phi).$
Example 3.17.
Let $G/H$ be cyclic with representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$. Suppose $\pi(hs^{i})=I_{m}$. Then,
$\rho(hs^{i})=\begin{bmatrix}0&\dots&&0&\pi(s^{-i})&0&\dots&\\ \vdots&&&&0&\pi(s^{-i})&0&\dots\\ &\ddots&&&\vdots&&\ddots&\\ &&&&&&&\pi(s^{-i})\\ \pi(s^{n-i})&&&&&&&0\\ 0&\pi(s^{n-i})&&&&&&\vdots\\ \vdots&0&\ddots&&&&&\\ &&0&\pi(s^{n-i})&&&&\\ \end{bmatrix}=\theta^{i}.$
Corollary 3.17.1.
$\tilde{L}\cong\tilde{\pi}(G)\iff ker(\tilde{\pi})\subset H$.
Proof.
[…]
By Lemma 3.16, $ker(\Phi)$ can be realized as a subgroup of $G/H$ and we denote this embedding by $\Omega:ker(\Phi)\xhookrightarrow{}G/H$. Let $G^{\prime}$ be the unique subgroup of $G$ such that $H\subset G^{\prime}\subset G$ and $ker(\Phi)=G^{\prime}/H$.
In fact, we have a more precise statement.
Lemma 3.18.
Let $H\subset G^{\prime}\subset G$. Then
$ker(\Phi)=G^{\prime}/H$ under $\Omega$ $\iff$ $G^{\prime}$ is the largest subgroup of $G$ such that $\pi(G^{\prime})=\pi(H)$.
In particular, $ker(\Phi)=G/H$ under $\Omega$ $\iff$ $\pi(G)=\pi(H)$.
Proof.
Now $H\subset G^{\prime}\subset G$. We have $G^{\prime}=\bigsqcup_{i\in I}(G^{\prime}\cap Hs_{i})$ where $I=\{i\in\{1,2,...,n\}\ |\ G^{\prime}\cap Hs_{i}\neq\emptyset\}$. Also $(G^{\prime}\cap Hs_{i})\neq\emptyset\iff s_{i}\in G^{\prime}$. Hence $I=\{i\in\{1,2,...,n\}\ |\ s_{i}\in G^{\prime}\}$. For $i\in I$, $(G^{\prime}\cap Hs_{i})=Hs_{i}$. Hence $G^{\prime}=\bigsqcup_{i\in I}Hs_{i}$ and $G^{\prime}/H=\{s_{i}H\}_{i\in I}$. Now the following argument completes the proof.
[…]
Theorem 3.19.
Suppose $\pi(G)=\pi(H)$. Then the exact sequence
(3.3) $1\rightarrow ker(\Phi)\xhookrightarrow{}L\rightarrow\pi(G)\rightarrow 1$
is right split and $L\cong\pi(G)\ltimes G/H$ for some semidirect product group law.
Proof.
Since $\pi(G)=\pi(H)$, $ker(\Phi)=G/H$ under $\Omega$. Consider $\iota:\pi(G)\rightarrow L$ given by $\iota(x)=(\pi(h),\rho(h))$ where $x=\pi(h)$ for some $h\in H$. If $x=\pi(h)=\pi(h^{\prime})$ for some $h,h^{\prime}\in H$ then $\pi(hh^{\prime-1})=I$. Hence by matrix calculation above, $\rho(hh^{\prime-1})=I$. Hence $\rho(h)=\rho(h^{\prime})$. Thus $\iota$ is well defined. Clearly, $\iota$ is a homomorphism and $\Phi\circ\iota=id_{\tilde{\pi}(G)}$, therefore the exact sequence splits.∎
Corollary 3.19.1.
Let $G/H$ be cyclic <sH><sH>. If $ker(\pi)\cap Hs\neq\emptyset$ then the exact sequence 3.3 is right split.
Proof.
Since $ker(\pi)\cap Hs\neq\emptyset$, there is an $h_{1}\in H$ such that $\pi(h_{1}s)=1$. Since $H$ is normal in $G$, $(h_{1}s)^{i}=h_{i}s^{i}$ for some $h_{i}\in H$. Hence $\pi(h_{i}s^{i})=\pi(h_{1}s)^{i}=1$. Hence for all $i$, $ker(\pi)\cap Hs^{i}\neq\emptyset$. Thus $ker(\Phi)\cong G/H$.∎
Corollary 3.19.2.
Let $\{\rho(s_{i})\}_{i}$ form a multiplicatively closed subset of $\rho(G)$ and let $(n,|\pi(G)|)=1$. Then the exact sequence 3.3 is right split.
Proof.
Since $|G/H|=n$, for any $i$, $(s_{i}H)^{n}=H$. Thus $\rho(s_{i})^{n}=I$. Hence $\pi(s_{i})^{n}=I$. Since $\pi(s_{i})\in\pi(G)$, we have $\pi(s_{i})^{|\pi(G)|}=I$. Now $(n,|\pi(G)|)=1$. Hence, $\pi(s_{i})=1$ for all $i$. Thus $\pi(H)=\pi(G)$. ∎
Recall the results from [8] that we discussed earlier in Sec. 3.2. In this case, we have
Corollary 3.19.3.
For $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\times\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$.
Proof.
Since $\tilde{\pi}(H)=\tilde{\pi}(G)$ as in Theorem 3.12, we have from above that ${\rm PSL}_{2}(\mathbb{F}_{p})\ltimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law as in 3.19). Since $Aut(\mathbb{Z}/2\mathbb{Z})$ is trivial, we have that semidirect product is indeed direct. ∎
We can generalize Lemma 3.16, Corollary 3.17.1, Lemma 3.18 and Theorem 3.19. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
Theorem 3.20.
Let $N=\{(\pi(h^{\prime}),\rho(h^{\prime}))|h^{\prime}\in H^{\prime}\}$ and $\tilde{N}=\{(\tilde{\pi}(h^{\prime}),\tilde{\rho}(h^{\prime}))|h^{\prime}\in H^{\prime}\}$.
  1. (1)
    The surjective projection $N\rightarrow\rho(H^{\prime})$ is an isomorphism.
  2. (2)
    Consider the surjective projection $\xi:N\rightarrow\pi(H^{\prime})$. We have $\omega:ker(\xi)\xhookrightarrow{}H^{\prime}/(H^{\prime}\cap H)$ and $ker(\xi)\cong(ker(\pi)\cap H^{\prime})/(ker(\pi)\cap H^{\prime}\cap H)$.
  3. (3)
    $\tilde{N}\cong\tilde{\pi}(H^{\prime})\iff ker(\tilde{\pi})\cap H^{\prime}\subset H$.
  4. (4)
    Let $(H^{\prime}\cap H)\subset G^{\prime}\subset H^{\prime}$ then
    $ker(\xi)=G^{\prime}/(H^{\prime}\cap H)$ under $\omega$ $\iff$ $G^{\prime}$ is largest subgroup of $H^{\prime}$ such that $\pi(G^{\prime})=\pi(H^{\prime}\cap H)$.
    In particular, $ker(\xi)=H^{\prime}/(H^{\prime}\cap H)$ under $\omega$ $\iff$ $\pi(H^{\prime})=\pi(H^{\prime}\cap H)$.
  5. (5)
    Suppose $\pi(H^{\prime})=\pi(H^{\prime}\cap H)$. Then the exact sequence
    (3.4) $1\rightarrow ker(\xi)\rightarrow N\rightarrow\pi(H^{\prime})\rightarrow 1$
    is right split and $N\cong\pi(H^{\prime})\ltimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law.
  6. (6)
    Let $H\subset H^{\prime}$. Suppose $\pi(G)=\pi(H)$. Then the exact sequence 3.4 is right split and $N\cong\pi(H^{\prime})\ltimes H^{\prime}/H$ for some semidirect product group law.
Acknowledgements: Both the authors would like to thank Prof. Joachim König for a prompt reply and useful discussion over email about the results by Arias-de-Reyna & König in [1]. The second author Shubham Jaiswal would like to acknowledge support of IISER Pune Institute Scholarship during this work.